OpenStudy (doreenhead1234):

Bryce deposited $100 in a savings account earning 5% interest, compounded annually. To the nearest cent, how much interest will he earn in 3 years?

563blackghost (563blackghost):

Use this equation. \(\huge\bf{A=P(1 + \frac{r}{n})^{n \times t}}\) `P = 100` `r = 0.05` `n = 1` `t = 3` \(\huge\bf{A=100(1 + \frac{0.05}{1})^{1 \times 3}}\) What you find you will plug it into this equation to find the interest. \(\huge\bf{A=P+I}\) `A = what you find` `P = 100`

563blackghost (563blackghost):

@doreenhead1234

likeabossssssss (likeabossssssss):

\(\Huge\color{#EB00FF}{\text{WELCOME}}\) \(\Huge\color{blue}{\text{TO}}\) \(\Huge\color{green}{\text{OPEN}}\)\(\Huge\color{purple }{\text{ STUDY!!!!!!!!!!!}}\) \(\Huge\heartsuit\)

OpenStudy (doreenhead1234):

Thank u likeabossssssss